# Correlation between the charge radii difference in mirror partner nuclei and the symmetry energy slope

### v1

@@ Abstract
Correlation between the charge radii difference of mirror partner nuclei $\Delta{R_{\mathrm{ch}}}$ and the slope parameter $L$ of symmetry energy has been built to ascertain the equation of state of isospin asymmetric nuclear matter. In this work, the influence of pairing correlations and isoscalar compression modulus on the $\Delta{R_{\mathrm{ch}}}$ are systematically investigated by Skyrme density functionals theory. The calculated results suggest that the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ is decreased by the surface pairing correlations. The slope parameter deduced from the difference of charge radii of mirror-pair nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe falls into the range of $L=42.57$-$50.64$ MeV, namely the rather soft equation of state of asymmetric nuclear matter. Besides, the range of slope parameter can also be influenced by the effective forces classified by various isoscalar incompressibility coefficients.
@@ I INTRODUCTION
Nuclear symmetry energy can be generally employed to comprehend the underlying physical mechanism from the terrestrial nuclei to dense astrophysical events Steiner et al. (2005); Li et al. (2008); Roca-Maza and Paar (2018); Ji et al. (2019); Zhang et al. (2020); Xu et al. (2022a). So far, plenty of methods are undertaken to constrain the nuclear symmetry energy, such as the extracted neutron skin thickness (NST) Oyamatsu et al. (1998); Alex Brown (2000); Furnstahl (2002); Liu et al. (2011); Zhang and Chen (2013); Xu et al. (2020), properties of giant resonances Cao and Ma (2008); Cao et al. (2015); Roca-Maza et al. (2015, 2016); Cheng et al. (2023), quantities of the heavy-ion collision Lynch et al. (2009); Toro et al. (2010); Kumar et al. (2011); Xie et al. (2013); Feng and Jin (2010); Yu et al. (2020), etc. However, a unified density-dependence of symmetry energy is hardly determined due to the uncertainties suffering from different theoretical models. Thus more alternative observables should be urgently required in constraining the equation of state (EoS) of asymmetric nuclear matter. The difference of charge radii of mirror-pair nuclei has been proposed as a tentative probe to evaluate the slope parameter of symmetry energy $L$ Wang and Li (2013); Brown (2017). And then the precisely measured charge radii differences in mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe are thereby used to validate the range of slope parameter $L$ Brown et al. (2020); Pineda et al. (2021); König et al. (2023).
@@
Charge radii differences in mirror pair nuclei are strongly associated with the neutron-skin thickness of neutron-rich nuclei and with the slope parameter $L$ Wang and Li (2013); Brown (2017). As shown in Ref. Yang and Piekarewicz (2018), the calculations performed by relativistic energy density functionals (EDFs) further inspect the validity of this assumption. It points out that the difference in proton radii of mirror nuclei can be indeed employed to constrain the EoS of isospin asymmetry nuclear matter. This may provide an alternative approach to extract the content information about the neutron skin thickness as well Ding et al. (2023). Recent study suggests that mirror-difference in nuclear charge radii is proportional to the isospin asymmetry $I=(N-Z)/A$, where $N$ and $Z$ represent the neutron and proton number of a nucleus, respectively, and $A=N+Z$ is the corresponding mass number Novario et al. (2023). Here the same scenario can also be observed between the difference of binding energy per nucleon of mirror nuclei and the Coulomb asymmetry. Furthermore, combing the accurately detected charge radii differences in mirror partner nuclei, the upper limit range of symmetry energy slope has been determined at $L\approx 100$ MeV Bano et al. (2023).
@@
Method mentioned above requires that more accurate charge radii data should be obtained through different nuclear models. With the improvement of experimental technique, high-precision charge radii of exotic nuclei can be possibly detected Yang et al. (2023a). Thus more charge radii data of nuclei far away from the $\beta$-stability line are compiled Angeli and Marinova (2013); Li et al. (2021). This provides a fundamental guideline for theoretical study to access the charge radii of mirror partner nuclei. A reliable description of nuclear charge radii is influenced by various mechanisms, such as shape deformation Lalazissis et al. (1996); Togashi et al. (2016); An et al. (2022a, 2023a), pairing correlations Reinhard and Nazarewicz (2017); An et al. (2020, 2022b), shell closure effects Nakada (2019); Perera et al. (2021), etc. As is well known, pairing correlations play an important role in describing the ground state properties of finite nuclei Meng (1998); Geng et al. (2003); An et al. (2018) or the quasiparticle resonant states Zhang and Qu (2020). Especially, it is mentioned that the radius of neutron-rich nuclei can also be influenced evidently by the pairing correlations Zhang et al. (2017).
@@
As demonstrated in Ref. Reinhard and Nazarewicz (2022), pairing correlations should be taken into account in calculating the nuclear charge radii. Shown in Ref. Huang et al. (2023), the density energy functionals consisting of Skyrme and covariant models are undertaken to systematically investigate the correlations between the charge radii differences in mirror partner nuclei and slope parameter $L$. The calculated results suggest the linear correlation between the mirror difference in charge radii and slope parameter $L$ is further decreased by considering the pairing effects. In contrast, the correlation between the neutron-skin thickness of a neutron-rich nucleus and the proton radii difference of the corresponding mirror nuclei is mostly enhanced by pairing effects Ding et al. (2023). Meanwhile, it is also mentioned that the compression modulus of symmetric nuclear matter has an influence on determining the slope parameter $L$ An et al. (2023b).
@@
The density-dependence of nuclear symmetry energy, which is mostly determined by the slope parameter $L$, is associated with the isovector-sensitive indicators in the EoS of isospin asymmetry systems. In Ref. Yoshida and Sagawa (2004), it suggests that the correlation between the isoscalar incompressibility $K$ and the NST is not clear. Further study demonstrates that the correlations between $K$ and the isovector parameters are generally weaker in comparison with the strong correlation between NST and symmetry energy coefficients Yoshida and Sagawa (2006); Xu et al. (2021). Actually, the nuclear matter properties recognizing the isoscalar and isovector counterparts are mutually correlated with the effective model parameters. This means the influence coming from the isoscalar sector is non-negligible in discussing the isovector properties Chatterjee et al. (2017); Mondal and Gulminelli (2023). As suggested in Refs. Piekarewicz (2004); Chen and Gu (2012); Kumar et al. (2021), the compression modulus of symmetric nuclear matter is sensitive to the density dependence of symmetry energy. Therefore, it is also essential to investigate the correlation between the difference of charge radii of mirror-partner nuclei and the slope parameter $L$ under various incompressibility coefficients $K$. In this work, the influence of pairing correlations and isoscalar compression modulus on the charge radii differences of mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe are investigated by the spherical Skyrme EDFs. Furthermore, the range of slope parameter $L$ deriving from the charge radii differences of mirror-pair nuclei is discussed as well.
@@
The structure of the paper is organized as the follows. Sec. II is devoted to presenting the theoretical framework briefly. In Sec. III, the numerical results and discussion are provided. Finally, a summary and outlook is given in Sec. IV.
@@ II THEORETICAL FRAMEWORK
The sophisticated Skyrme-type energy density functional, which was expressed as an effective zero-range force between nucleons with density-dependent and momentum-dependent terms, has made remarkable success in describing various physical phenomena Bender et al. (2003); Pei et al. (2008, 2013); Wang et al. (2016); Zhang et al. (2017); Jiang and Wang (2018); Qu and Zhang (2019); Wu and Guo (2019); Jiang and Wang (2020a, b); Zhang and Qu (2020); Pei et al. (2014); Chai et al. (2020, 2020); Zhang et al. (2021); Chai et al. (2022); Wu et al. (2022); Huo et al. (2023). In general, the effective force sets have been determined by calibrating the properties of finite nuclei and symmetry nuclear matter at saturation density $\rho_{0}$($\approx 0.16~{}\mathrm{fm}^{-3}$). It is notably mentioned that the Skyrme EDF can give analytical expression of all variables characterizing infinite nuclear matter (see, e.g., Nakada (2003); Dutra et al. (2012) for details). In this work, the Skyrme-like effective interaction has been recalled as follows Köhler (1976); Chabanat et al. (1997, 1998),
@@
Here, $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ and $\mathbf{R}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$ are related to the positions of two nucleons $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$, $\mathbf{P}=(\nabla_{1}-\nabla_{2})/2\mathrm{i}$ is the relative momentum operator and $\mathbf{P^{\prime}}$ is its complex conjugate acting on the left, and $\mathbf{P_{\sigma}}=(1+\vec{\sigma}_{1}\cdot\vec{\sigma}_{2})/2$ is the spin exchange operator that controls the relative strength of the $S=0$ and $S=1$ channels for a given term in the two-body interactions, with $\vec{\sigma}_{1(2)}$ being the Pauli matrices. The last term represents the spin-orbit force where $\sigma=\vec{\sigma}_{1}+\vec{\sigma}_{2}$. The quantities $\alpha$, $t_{i}$ and $x_{i}$ ($i=0$-3) represent the parameters of the Skyrme forces used in this work.
Solving the Schrödinger-like equations by the self-consistent iteration leads to the eigenenergies and wave functions of constituent nucleons. The pairing correlations can be generally treated either by the BCS method or by the Bogoliubov transformation Meng and Ring (1996); Meng (1998); Geng et al. (2003). In this work, spherical Skyrme Hartree-Fock-Bogoliubov (HFB) approach is utilized for all calculations Bennaceur and Dobaczewski (2005). The density-dependent zero-range pairing force is employed as follows Cao et al. (2012),
@@
where $\rho(\mathbf{r})$ is the baryon density in coordinate space and $\rho_{0}=0.16~{}\mathrm{fm}^{-3}$ represents the nuclear saturation density. The value of $\eta$ is taken as $0.0$, $0.5$, or $1.0$ for volume-, mixed-, or surface-type pairing interactions, respectively. The quantity $V_{0}$ is adjusted by calibrating the empirical pairing gaps with three-point formula Cao et al. (2012); An et al. (2018), and the values of $V_{0}$ are 188.2 MeV fm3, 370.2 MeV fm3 and 509.6 MeV fm3 for the corresponding volume-, mixed-, and surface-type pairing interactions, respectively.
After achieving the convergence of the total binding energy, the root-mean square (rms) radii of the neutron and proton matter, and the neutron-skin thickness can be obtained naturally. The quantity of nuclear charge radius ($R_{\mathrm{ch}}$) can be calculated through (in units of fm2)
@@
Table 1: Parameters of the forces used in this work, where the parameter $\alpha=1/6$. Saturation properties of the Skyrme parametrization sets used in this work, such as the incompressibility coefficients $K$ (MeV), slope parameter $L$ (MeV) and symmetry energy $E_{\mathrm{sys}}$ (MeV) at saturation density $\rho_{0}$ (fm-3), are also shown.
   $K$    Sets     $t_{0}$      $t_{1}$     $t_{2}$     $t_{3}$     $x_{0}$     $x_{1}$     $x_{2}$     $x_{3}$     $W_{0}$     $L$     $E_{\mathrm{sys}}$
@@
The density dependence of symmetry energy can be expanded around the saturation density $\rho_{0}$ as follows
$E_{\mathrm{sys}}(\rho)\approx{E_{\mathrm{sys}}(\rho_{0})}+\frac{L}{3}\left(\frac{\rho-\rho_{0}}{\rho_{0}}\right)+\frac{K_{\mathrm{sym}}}{18}\left(\frac{\rho-\rho_{0}}{\rho_{0}}\right)^{2}+\dots,$ (4)
@@
$L$ $=$ $3\rho_{0}\frac{\partial{E_{\mathrm{sys}}(\rho)}}{{{\partial\rho}}}\Large{|}_{\rho=\rho_{0}},$ (5)
$K_{\mathrm{sym}}$ $=$ $9\rho_{0}^{2}\frac{\partial^{2}E_{\mathrm{sys}}(\rho)}{{\partial\rho^{2}}}\Large{|}_{\rho=\rho_{0}}.$ (6)
Also, the isoscalar incompressibility $K$ of symmetric nuclear matter is recalled as
@@
Here, the quantity $E_{0}(\rho)$ is the energy per particle of symmetric nuclear matter Ding et al. (2023). The nuclear incompressibility can be extracted from the measurements of the isoscalar giant monopole resonance (ISGMR) in medium-heavy nuclei Colò et al. (2004); Cao et al. (2012). In addition, it is also mentioned that the incompressibility of symmetric nuclear matter can be deduced from the $\alpha$-decay properties Seif (2006). In Table 1, parameters of the Skyrme forces used in this work are shown. The corresponding saturation properties in infinite nuclear matter derived from the Skyrme forces, such as isoscalar incompressibility $K$, slope parameter $L$, and symmetry energy $E_{\mathrm{sys}}$, are also shown explicitly. Both the slope parameter $L$ and symmetry energy $E_{\mathrm{sys}}$ at saturation density cover a wide range.
@@
Linear correlation between the charge radii differences in the mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe and the slope parameter of symmetry energy has been sequentially employed to constrain the equation of state of isospin asymmetric nuclear matter Brown et al. (2020); Pineda et al. (2021); König et al. (2023).
Table 2: $R_{\mathrm{ch}}$ and $\Delta{R_{\mathrm{ch}}}$ database for the $A=32$, $36$, $38$ and $54$ mirror-pair nuclei. The parentheses on the values of charge radii and the difference of charge radii are shown with systematic uncertainties Angeli and Marinova (2013); Miller et al. (2019); Pineda et al. (2021); Brown et al. (2020); König et al. (2023).
@@
In Table 2, the experimental data of $R_{\mathrm{ch}}$ and $\Delta{R_{\mathrm{ch}}}$ for $A=32$, $36$, $38$ and $54$ mirror-pair nuclei are given. As mentioned above, nuclear charge radius is influenced by the pairing correlations Reinhard and Nazarewicz (2017); An et al. (2020, 2022b). Thus the influence on the correlations between the charge radii difference in mirror partner nuclei and the slope parameter $L$ should be systematically investigated.
@@
Figure 1: (Color online)Charge radii difference $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 54Ni-54Fe as a function of slope parameter $L$ is depicted under various pairing forces, e.g., no pairing (a), volume type $\eta=0.0$ (b), mixed type $\eta=0.5$ (c) and surface type $\eta=1.0$ (d) pairing forces. The Pearson coefficients $R$ are given in the parenthesis. The database takes from Refs. Angeli and Marinova (2013); Li et al. (2021); Pineda et al. (2021) and the symmetric error bands are labeled by light blue.
@@
In Fig. 1, plots for charge radii difference $\Delta{R_{\mathrm{ch}}}$ in 54Ni-54Fe against symmetry energy slope $L$ are depicted by various pairing forces. For the convenience of discussion, the Pearson coefficient $R$, which is used to measure the degree of correlation between two variables, is employed in our study. Without considering the pairing correlations, the Pearson coefficient $R$ between $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 54Ni-54Fe and $L$ is given to be $0.993$. The same scenario can also be found in mixed-type pairing forces. For the volume-type pairing force, the Pearson coefficient $R$ is almost equivalent to no pairing and $\eta=0.0$ cases. As demonstrated in Ref. Huang et al. (2023), the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ is decreased by the pairing effects. From Fig. 1, one can find that the surface-type pairing force actually decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ in mirror pair nuclei 54Ni-54Fe, in which the Pearson coefficient $R$ falls into $0.882$. Actually, more particles are scattered into the higher single levels under the surface pairing force. This leads to the rather poor correlation between the difference of charge radii of mirror partner nuclei and the slope parameter $L$.
@@
Figure 2: (Color online)Charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar as a function of slope parameter $L$ are depicted under no pairing and mixed-type $\eta=0.5$ pairing forces. The Pearson coefficients $R$ are given in the parenthesis. The database takes from Refs. Angeli and Marinova (2013); Li et al. (2021); Brown et al. (2020); König et al. (2023) and the symmetric error bands are labeled by light blue.

### v2

@@ Abstract
A correlation between the charge radii difference of mirror partner nuclei $\Delta{R_{\mathrm{ch}}}$ and the slope parameter $L$ of symmetry energy has been built to ascertain the equation of state of isospin asymmetric nuclear matter. In this work, the influences of pairing correlations and isoscalar compression modulus on the $\Delta{R_{\mathrm{ch}}}$ are systematically investigated based on the Skyrme energy density functional theory. The calculated results suggest that the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ is decreased by the surface pairing correlations. The slope parameter deduced from the difference of charge radii of mirror-pair nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe falls into the range of $L=42.57$-$50.64$ MeV, that is, the rather soft equation of state of asymmetric nuclear matter. Besides, the range of the slope parameter can also be influenced by the effective forces classified by various isoscalar incompressibility coefficients.
@@ I INTRODUCTION
Nuclear symmetry energy is generally employed to comprehend the underlying physical mechanism from terrestrial nuclei to dense astrophysical events Steiner et al. (2005); Li et al. (2008); Roca-Maza and Paar (2018); Ji et al. (2019); Zhang et al. (2020); Xu et al. (2022a). So far, plenty of methods have been applied to constrain the nuclear symmetry energy, such as the extracted neutron skin thickness (NST) Oyamatsu et al. (1998); Alex Brown (2000); Furnstahl (2002); Liu et al. (2011); Zhang and Chen (2013); Xu et al. (2020), properties of giant resonances Cao and Ma (2008); Cao et al. (2015); Roca-Maza et al. (2015, 2016); Cheng et al. (2023), and quantities of the heavy-ion collision Lynch et al. (2009); Toro et al. (2010); Kumar et al. (2011); Xie et al. (2013); Feng and Jin (2010); Yu et al. (2020). However, a unified density-dependence of symmetry energy is hardly determined owing to the uncertainties of the different theoretical models. Thus, alternative observables are urgently required for constraining the equation of state (EoS) of asymmetric nuclear matter. The difference in charge radii of mirror-pair nuclei has been proposed as a tentative probe to evaluate the slope parameter of symmetry energy $L$ Wang and Li (2013); Brown (2017). Therefore, the precisely measured charge radii differences in mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe have been thereby used to validate the range of the slope parameter $L$ Brown et al. (2020); Pineda et al. (2021); König et al. (2024).
@@
Charge radii differences in mirror pair nuclei are strongly associated with the neutron-skin thickness of neutron-rich nuclei and with the slope parameter $L$ Wang and Li (2013); Brown (2017). As shown in Ref. Yang and Piekarewicz (2018), the calculations performed using relativistic energy density functionals (EDFs) further inspect the validity of this assumption. It points out that the difference in proton radii of mirror nuclei can be indeed employed to constrain the EoS of isospin asymmetry nuclear matter. This may provide an alternative approach to extract content information about the neutron skin thickness Ding et al. (2023). A recent study suggests that the mirror-difference in nuclear charge radii is proportional to the isospin asymmetry $I=(N-Z)/A$, where $N$ and $Z$ represent the neutron and proton number of a nucleus, respectively, and $A=N+Z$ is the corresponding mass number Novario et al. (2023). Here, the same scenario can also be observed between the difference in binding energy per nucleon of mirror nuclei and the Coulomb asymmetry. Furthermore, combing the accurately detected charge radii differences in mirror partner nuclei, the upper limit range of the symmetry energy slope has been determined to be $L\approx 100$ MeV Bano et al. (2023).
@@
This method requires more accurate charge radii data, which should be obtained through different nuclear models. With the improvements in the experimental technique, high-precision charge radii of exotic nuclei can be detected Yang et al. (2023a). Thus, more charge radii data of nuclei far away from the $\beta$-stability line can be compiled Angeli and Marinova (2013); Li et al. (2021). This provides a fundamental guideline for theoretical studies to access the charge radii of mirror partner nuclei. A reliable description of nuclear charge radii is influenced by various mechanisms, such as shape deformation Lalazissis et al. (1996); Togashi et al. (2016); An et al. (2022a, 2023a), pairing correlations Reinhard and Nazarewicz (2017); An et al. (2020, 2022b), and shell closure effects Nakada (2019); Perera et al. (2021). As is well known, pairing correlations play an important role in describing the ground state properties of finite nuclei Meng (1998); Geng et al. (2003); An et al. (2018) or the quasiparticle resonant states Zhang and Qu (2020). In particular, it has been mentioned that the radius of neutron-rich nuclei can also be influenced by the pairing correlations Zhang et al. (2017).
@@
As demonstrated in Ref. Reinhard and Nazarewicz (2022), the pairing correlations should be taken into account in calculating the nuclear charge radii. As presented in Ref. Huang et al. (2023), the density energy functionals consisting of Skyrme and covariant models are used to systematically investigate the correlations between the charge radii differences in mirror partner nuclei and slope parameter $L$. The calculated results suggest that the linear correlation between the mirror difference in charge radii and slope parameter $L$ is further decreased when considering the pairing effects. In contrast, the correlation between the neutron-skin thickness of a neutron-rich nucleus and the proton radii difference of the corresponding mirror nuclei is mostly enhanced by pairing effects Ding et al. (2023). Meanwhile, it has also been mentioned that the compression modulus of symmetric nuclear matter has an influence on determining the slope parameter $L$ An et al. (2023b).
@@
The density-dependence of nuclear symmetry energy, which is mostly determined by the slope parameter $L$, is associated with the isovector-sensitive indicators in the EoS of isospin asymmetry systems. In Ref. Yoshida and Sagawa (2004), it is stated that the correlation between the isoscalar incompressibility $K$ and the NST is not clear. Further research demonstrated that the correlations between $K$ and the isovector parameters are generally weaker in comparison with the strong correlation between NST and symmetry energy coefficients Yoshida and Sagawa (2006); Xu et al. (2021). Actually, the nuclear matter properties recognizing the isoscalar and isovector counterparts are mutually correlated with the effective model parameters. This means that the influence coming from the isoscalar sector is non-negligible in discussing the isovector properties Chatterjee et al. (2017); Mondal and Gulminelli (2023). As suggested in Refs. Piekarewicz (2004); Chen and Gu (2012); Kumar et al. (2021), the compression modulus of symmetric nuclear matter is sensitive to the density dependence of symmetry energy. Therefore, it is also essential to investigate the correlation between the difference in charge radii of mirror-partner nuclei and the slope parameter $L$ under various incompressibility coefficients $K$. In this work, the influences of pairing correlations and isoscalar compression modulus on the charge radii differences of mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe are investigated using the spherical Skyrme EDFs. Furthermore, the range of the slope parameter $L$ derived from the charge radii differences of mirror-pair nuclei is discussed.
@@
The contents of this paper are organized as follows. Sec. II summarizes the theoretical framework. In Sec. III, the numerical results and discussion are provided. Finally, a summary and outlook are presented in Sec. IV.
@@ II THEORETICAL FRAMEWORK
The sophisticated Skyrme-type EDF, which is expressed as an effective zero-range force between nucleons with density-dependent and momentum-dependent terms, has achieved remarkable success in describing various physical phenomena Bender et al. (2003); Pei et al. (2008, 2013); Wang et al. (2016); Zhang et al. (2017); Jiang and Wang (2018); Qu and Zhang (2019); Wu and Guo (2019); Jiang and Wang (2020a, b); Zhang and Qu (2020); Pei et al. (2014); Chai et al. (2020, 2020); Zhang et al. (2021); Chai et al. (2022); Wu et al. (2022); Huo et al. (2023). In general, the effective force sets are determined by calibrating the properties of the finite nuclei and symmetric nuclear matter at saturation density $\rho_{0}$($\approx 0.16~{}\mathrm{fm}^{-3}$). Notably, it has been mentioned that some interactions can be used to characterize the bulk properties of finite nuclei and infinite nuclear matter (see, e.g., Nakada (2003); Dutra et al. (2012) for details). In this work, the Skyrme-like effective interaction has been recalled as follows Köhler (1976); Chabanat et al. (1997, 1998):
@@
Here, $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ and $\mathbf{R}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$ are related to the positions of two nucleons $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$; $\mathbf{P}=(\nabla_{1}-\nabla_{2})/2\mathrm{i}$ is the relative momentum operator and $\mathbf{P^{\prime}}$ is its complex conjugate acting on the left; and $\mathbf{P_{\sigma}}=(1+\vec{\sigma}_{1}\cdot\vec{\sigma}_{2})/2$ is the spin exchange operator that controls the relative strength of the $S=0$ and $S=1$ channels for a given term in the two-body interactions, with $\vec{\sigma}_{1(2)}$ being the Pauli matrices. The last term represents the spin-orbit force where $\sigma=\vec{\sigma}_{1}+\vec{\sigma}_{2}$. The quantities $\alpha$, $t_{i}$, and $x_{i}$ ($i=0$-3) represent the parameters of the Skyrme forces used in this work.
Solving the Schrödinger-like equations by self-consistent iteration leads to the eigenenergies and wave functions of the constituent nucleons. The pairing correlations can be generally treated either by the BCS method or by the Bogoliubov transformation Meng and Ring (1996); Meng (1998); Geng et al. (2003). In this work, the spherical Skyrme Hartree-Fock-Bogoliubov (HFB) approach is utilized for all the calculations Bennaceur and Dobaczewski (2005). The density-dependent zero-range pairing force is employed as follows Cao et al. (2012):
@@
where $\rho(\mathbf{r})$ is the baryon density in the coordinate space and $\rho_{0}=0.16~{}\mathrm{fm}^{-3}$ represents the nuclear saturation density. The value of $\eta$ is taken as $0.0$, $0.5$, or $1.0$ for volume-, mixed-, or surface-type pairing interactions, respectively. The quantity $V_{0}$ is adjusted by calibrating the empirical pairing gaps with the three-point formula Cao et al. (2012); An et al. (2018), and the values of $V_{0}$ are 188.2, 370.2, and 509.6 MeV fm3 for the corresponding volume-, mixed-, and surface-type pairing interactions, respectively.
After achieving the convergence of the total binding energy, the root-mean square (rms) radii of the neutron and proton matter, and the neutron-skin thickness can be obtained naturally. The quantity of nuclear charge radius ($R_{\mathrm{ch}}$) can be calculated as follows (in units of fm2):
@@
Table 1: Parameters of the forces used in this work, where the parameter $\alpha=1/6$. Saturation properties of the Skyrme parametrization sets used in this work, such as the incompressibility coefficients $K$ (MeV), slope parameter $L$ (MeV), and symmetry energy $E_{\mathrm{sym}}$ (MeV) at saturation density $\rho_{0}$ (fm-3), are also listed.
   $K$    Sets     $t_{0}$      $t_{1}$     $t_{2}$     $t_{3}$     $x_{0}$     $x_{1}$     $x_{2}$     $x_{3}$     $W_{0}$     $L$     $E_{\mathrm{sym}}$
@@
The density dependence of symmetry energy can be expanded around the saturation density $\rho_{0}$ as follows:
$E_{\mathrm{sym}}(\rho)\approx{E_{\mathrm{sym}}}(\rho_{0})+\frac{L}{3}\left(\frac{\rho-\rho_{0}}{\rho_{0}}\right)+\frac{K_{\mathrm{sym}}}{18}\left(\frac{\rho-\rho_{0}}{\rho_{0}}\right)^{2}+\dots,$ (4)
@@
$L$ $=$ $3\rho_{0}\frac{\partial{E_{\mathrm{sym}}(\rho)}}{{{\partial\rho}}}\Large{|}_{\rho=\rho_{0}},$ (5)
$K_{\mathrm{sym}}$ $=$ $9\rho_{0}^{2}\frac{\partial^{2}E_{\mathrm{sym}}(\rho)}{{\partial\rho^{2}}}\Large{|}_{\rho=\rho_{0}}.$ (6)
Further, the isoscalar incompressibility $K$ of symmetric nuclear matter is recalled as
@@
Here, the quantity $E_{0}(\rho)$ is the energy per particle of symmetric nuclear matter Ding et al. (2023). The nuclear incompressibility can be extracted from the measurements of the isoscalar giant monopole resonance (ISGMR) in medium-heavy nuclei Colò et al. (2004); Cao et al. (2012). In addition, it is also mentioned that the incompressibility of symmetric nuclear matter can be deduced from the $\alpha$-decay properties Seif (2006). Table 1, presents the parameters of the Skyrme forces used in this work. The corresponding saturation properties in infinite nuclear matter derived from the Skyrme forces, such as isoscalar incompressibility $K$, slope parameter $L$, and symmetry energy $E_{\mathrm{sym}}$, are also listed explicitly. Both the slope parameter $L$ and symmetry energy $E_{\mathrm{sym}}$ at saturation density cover a wide range.
@@
The linear correlation between the charge radii differences in the mirror partner nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe and the slope parameter of symmetry energy have been sequentially employed to constrain the EoS of isospin asymmetric nuclear matter Brown et al. (2020); Pineda et al. (2021); König et al. (2024).
Table 2: $R_{\mathrm{ch}}$ and $\Delta{R_{\mathrm{ch}}}$ database for the $A=32$, $36$, $38$, and $54$ mirror-pair nuclei. The parentheses on the values of charge radii and the difference of charge radii show the systematic uncertainties Angeli and Marinova (2013); Miller et al. (2019); Pineda et al. (2021); Brown et al. (2020); König et al. (2024).
@@
In Table 2, the experimental data of $R_{\mathrm{ch}}$ and $\Delta{R_{\mathrm{ch}}}$ for $A=32$, $36$, $38$ and $54$ mirror-pair nuclei are given. As mentioned above, the nuclear charge radius is influenced by the pairing correlations Reinhard and Nazarewicz (2017); An et al. (2020, 2022b). Thus, the influence on the correlations between the charge radii difference in mirror partner nuclei and the slope parameter $L$ should be systematically investigated.
@@
Figure 1: (Color online) Charge radii difference $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 54Ni-54Fe as a function of slope parameter $L$ is depicted under various pairing forces: no pairing (a), volume type $\eta=0.0$ (b), mixed type $\eta=0.5$ (c), and surface type $\eta=1.0$ (d) pairing forces. The Pearson coefficients $R$ are presented in parentheses. The database is taken from Refs. Angeli and Marinova (2013); Li et al. (2021); Pineda et al. (2021), and the systematic error bands are represented in light blue.
@@
In Fig. 1, plots for the charge radii difference $\Delta{R_{\mathrm{ch}}}$ in 54Ni-54Fe against symmetry energy slope $L$ are shown for various pairing forces. For the convenience of discussion, the Pearson coefficient $R$, which is used to measure the degree of correlation between two variables, is employed in our study. Without considering the pairing correlations, the Pearson coefficient $R$ between the $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 54Ni-54Fe and $L$ is determined to be $0.993$. The same scenario can also be found in mixed-type pairing forces. For the volume-type pairing force, the Pearson coefficient $R$ is almost equivalent to those for the no pairing and $\eta=0.0$ cases. As demonstrated in Ref. Huang et al. (2023), the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ is decreased by the pairing effects. From Fig. 1, one can find that the surface-type pairing force actually decreases the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ in mirror pair nuclei 54Ni-54Fe, where the Pearson coefficient $R$ falls to $0.882$. Actually, more particles are scattered into the higher single levels under the surface pairing force. This leads to the rather poor correlation between the difference of charge radii of mirror partner nuclei and the slope parameter $L$.
@@
Figure 2: (Color online) Charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar as a function of slope parameter $L$ depicted under no pairing and mixed-type $\eta=0.5$ pairing forces. The Pearson coefficients $R$ are presented in parentheses. The database is taken from Refs. Angeli and Marinova (2013); Li et al. (2021); Brown et al. (2020); König et al. (2024), and the systematic error bands are represented in light blue.